the longer side of a rectangle is four times the length of the shorter side. if the diagonal is 17cm long, find the dimensions of the triangle
Answer by mananth(16949)

shorter side =x
longer side = 4x
diagonal = 17
they form a right triangle
17^2=x^2+(4x)^2
17^2=x^2+16x^2
17^2=17x^2
x^2= 17
x= +/- sqrt17
ignore negative value
x= root 17 shorter side
longer side = 4sqrt17